# Program evaluation with remotely sensed outcomes

### v2

@@ D.1 Algorithm details
Focusing on the numerator, if $n^{1/2}\mathbb{E}_{n}(\hat{U}\hat{Z})-n^{1/2}\mathbb{E}_{n}(\hat{U}Z)=o_{p}(1)$ and $n^{1/2}\mathbb{E}_{n}(\hat{U}Z)\rightsquigarrow\mathcal{N}(0,V)$, then $n^{1/2}\mathbb{E}_{n}(U\hat{Z})\rightsquigarrow\mathcal{N}(0,V)$ by Slutsky’s theorem.
[…]
Overall, we conclude that $n^{1/2}(\hat{\theta}-\theta)\rightsquigarrow\mathcal{N}\left[0,\frac{V}{\{\mathbb{E}(XZ)\}^{2}}\right]$ by Slutsky’s theorem. In the following lemmas, we prove the remaining high probability statements. ∎
@@
Under Proposition 4.2’s conditions, $n^{1/2}\mathbb{E}_{n}(\hat{U}\hat{Z})-n^{1/2}\mathbb{E}_{n}(\hat{U}Z)=o_{p}(1)$.
@@
Under Proposition 4.2’s conditions, $n^{1/2}\mathbb{E}_{n}(\hat{U}Z)\rightsquigarrow\mathcal{N}(0,V)$ where $V=v^{\top}\Sigma v$, $\Sigma_{ij}=cov(B_{i},B_{j})$, and
@@
We make three observations. First, if $\mathbb{E}(Z^{2})<\infty$, then by the Lindeberg-Levy central limit theorem, $n^{1/2}\{\mathbb{E}_{n}(B)-\mathbb{E}(B)\}\rightsquigarrow\mathcal{N}(0,\Sigma).$
@@
$=\mathbb{E}\{(\Delta^{e}-\theta\Delta^{o})Z\}=\mathbb{E}[\mathbb{E}\{(\Delta^{e}-\theta\Delta^{o})|Z\}]=0.$
@@
$n^{1/2}\mathbb{E}_{n}(\hat{U}Z)=n^{1/2}[h\{\mathbb{E}_{n}(B)\}-h\{\mathbb{E}(B)\}]\rightsquigarrow\mathcal{N}(0,[\nabla h\{\mathbb{E}(B)\}]^{\top}\Sigma[\nabla h\{\mathbb{E}(B)\}]).\qed$
@@
Under Proposition 4.2’s conditions, $\mathbb{E}_{n}(\hat{X}\hat{Z})-\mathbb{E}_{n}(\hat{X}Z)=o_{p}(1)$.
@@
[…]
$\mathbb{E}_{n}\{(\hat{X}-X)^{2}\}=\frac{1}{L}\sum_{\ell=1}^{L}\frac{1}{n_{\ell}}\sum_{i\in I_{\ell}}(\hat{X}_{i}-X_{i})^{2}=\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{E}_{\ell}\{(\hat{X}_{i}-X_{i})^{2}\},\quad\mathbb{E}_{\ell}(\cdot)=.\frac{1}{n_{\ell}}\sum_{i\in I_{\ell}}(\cdot).$
@@
[…]
@@
Under Proposition 4.2’s conditions, $\mathbb{E}_{n}(XZ)=\mathbb{E}(XZ)+o_{p}(1)$.
@@ Appendix E Discrete outcomes
Our identification results directly extend from $\mathcal{Y}=\{0,1\}$ to $\mathcal{Y}=\{y_{1},...,y_{|\mathcal{Y}|}\}$.
@@
Suppose Assumptions 2.1 and 2.2 hold.
  1. 1.
    If Assumption 2.3(i) holds, then
    $\Pr(R=r\mid D=d,S=e,X)=\sum_{y\in\mathcal{Y}}\Pr(R=r\mid Y=y,D=d,S=o)\Pr\{Y(d)=y\mid S=e,X\}.$
  2. 2.
    If Assumption 2.3(ii) holds, then
    $\Pr(R=r\mid D=d,S=e,X)=\sum_{y\in\mathcal{Y}}\Pr(R=r\mid Y=y,S=o)\Pr\{Y(d)=y\mid S=e,X\}.$
@@
By iterated expectations, we write
$\Pr(R\mid D=d,S=e,X)$ $=\sum_{y\in\mathcal{Y}}\Pr(R\mid Y=y,D=d,S=e,X)\Pr(Y=y\mid D=d,S=e,X)$
$=\sum_{y\in\mathcal{Y}}\Pr(R\mid Y=y,D=d,S=e,X)\Pr\{Y(d)=y\mid S=e,X\},$
[…]
@@
$\gamma_{y}(X,R)=\frac{\Pr(Y=y,S=o\mid X,R)}{\Pr(Y=y,S=o\mid X)}\mbox{ or }\gamma_{y,d}(X,R)=\frac{\Pr(Y=y,D=d,S=o\mid X,R)}{\Pr(Y=y,D=d,S=o\mid X)}.$
@@
Suppose Assumptions 2.1 and 2.2 hold.
@@
    If Assumption 2.3(i) holds, then $\pi_{d}(X,R)=\sum_{y\in\mathcal{Y}}\gamma_{y,d}(X,R)\Pr\{Y(d)=y\mid S=e,X\}.$

### v3

@@ D.4 Proof of Proposition 3
Focusing on the numerator, if $n^{1/2}\mathbb{E}_{n}(\hat{U}\hat{Z})=n^{1/2}\mathbb{E}_{n}(\hat{U}Z)+o_{p}(1)$ and $n^{1/2}\mathbb{E}_{n}(\hat{U}Z)\rightsquigarrow\mathcal{N}(0,V)$, then $n^{1/2}\mathbb{E}_{n}(U\hat{Z})\rightsquigarrow\mathcal{N}(0,V)$ by Slutsky’s theorem.
Focusing on the denominator, if $\mathbb{E}_{n}(\hat{X}\hat{Z})=\mathbb{E}_{n}(\hat{X}Z)+o_{p}(1)$, $\mathbb{E}_{n}(\hat{X}Z)=\mathbb{E}_{n}(XZ)+o_{p}(1)$, and $\mathbb{E}_{n}(XZ)=\mathbb{E}(XZ)+o_{p}(1)$, then $\mathbb{E}_{n}(\hat{X}\hat{Z})=\mathbb{E}(XZ)+o_{p}(1)$ by the continuous mapping theorem.
Overall, we conclude that $n^{1/2}(\hat{\theta}-\theta)\rightsquigarrow\mathcal{N}\left[0,\frac{V}{\{\mathbb{E}(XZ)\}^{2}}\right]$ by Slutsky’s theorem. In the following lemmas, we prove the remaining high probability statements and derive $V$. ∎
@@
Under Proposition 3’s conditions, $n^{1/2}\mathbb{E}_{n}(\hat{U}\hat{Z})=n^{1/2}\mathbb{E}_{n}(\hat{U}Z)+o_{p}(1)$.
@@
Under Proposition 3’s conditions, $n^{1/2}\mathbb{E}_{n}(\hat{U}Z)\rightsquigarrow\mathcal{N}(0,V)$ where $V=v^{\top}\Sigma v$, $\Sigma_{ij}=cov(B_{i},B_{j})$, and
@@
We make three observations. First, $\mathbb{E}(Z^{2})<\infty$ by Assumption 4, so by the Lindeberg-Levy central limit theorem, $n^{1/2}\{\mathbb{E}_{n}(B)-\mathbb{E}(B)\}\rightsquigarrow\mathcal{N}(0,\Sigma).$
@@
$=\mathbb{E}\{(\Delta^{e}-\theta\Delta^{o})Z\}=\mathbb{E}[\mathbb{E}\{(\Delta^{e}-\theta\Delta^{o})|Z\}Z]=0.$
@@
$n^{1/2}\mathbb{E}_{n}(\hat{U}Z)=n^{1/2}[h\{\mathbb{E}_{n}(B)\}-h\{\mathbb{E}(B)\}]\rightsquigarrow\mathcal{N}(0,[\nabla h\{\mathbb{E}(B)\}]^{\top}\Sigma[\nabla h\{\mathbb{E}(B)\}]).$
@@
Under Proposition 3’s conditions, $\mathbb{E}_{n}(\hat{X}\hat{Z})=\mathbb{E}_{n}(\hat{X}Z)+o_{p}(1)$.
@@
Since $\mathbb{E}_{n}(Z^{2})=\mathbb{E}(Z^{2})+o_{p}(1)$ when $\mathbb{E}(Z^{2})<\infty$ by the weak law of large numbers, it suffices to study
$\mathbb{E}_{n}\{(\hat{X}-X)^{2}\}=\frac{1}{L}\sum_{\ell=1}^{L}\frac{1}{n_{\ell}}\sum_{i\in I_{\ell}}(\hat{X}_{i}-X_{i})^{2}=\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{E}_{\ell}\{(\hat{X}_{i}-X_{i})^{2}\},\quad\mathbb{E}_{\ell}(\cdot)=\frac{1}{n_{\ell}}\sum_{i\in I_{\ell}}(\cdot).$
@@
For all $\delta>0$, this quantities vanishes for large $n_{\ell}=n/L$, so $\mathbb{E}_{\ell}\{(\hat{X}-X)^{2}\}=o_{p}(1)$. The continuous mapping theorem implies the desired result. ∎
@@
Under Proposition 3’s conditions, $\mathbb{E}_{n}(XZ)=\mathbb{E}(XZ)+o_{p}(1)$.
@@ Appendix E Discrete Outcomes
In this section, we extend our identification results and estimation strategy from $\mathcal{Y}=\{0,1\}$ to $\mathcal{Y}=\{y_{1},...,y_{K}\}$ for $K>2$. Our results in the main text follow from the special case with $K=2$ with $y_{1}=1$ and $y_{2}=0$.
E.1 Identification
As in the main text, our identification result allows for the covariates $X$ to be either discrete or continuous. We will also assume that the support of $R$ has at least $K$ elements (i.e., it is either multi-valued or continuous).
@@ E.1 Identification
Suppose Assumptions 1 and 2 hold.
  1. i.
    If Assumption 3(i) holds, then, for any $d\in\{0,1\}$, $x\in\mathcal{X}$ and $r\in\mathcal{R}$,
    $f_{R}(r\mid S=e,X,D=d)=\sum_{y\in\mathcal{Y}}f_{R}(r\mid S=o,X,D=d,Y=y)\Pr\{Y(d)=y\mid S=e,X\}.$
  2. ii.
    If Assumption 3(ii) holds, then, for any $d\in\{0,1\}$, $x\in\mathcal{X}$ and $r\in\mathcal{R}$,
    $f_{R}(r\mid S=e,X,D=d)=\sum_{y\in\mathcal{Y}}f_{R}(r\mid S=o,X,Y=y)\Pr\{Y(d)=y\mid S=e,X\}.$
@@
The former result specializes the general equation (5). The latter result specializes the general equation (3). ∎
@@
$\gamma_{y,d}(X,R)=\frac{\Pr(Y=y,D=d,S=o\mid X,R)}{\Pr(Y=y,D=d,S=o\mid X)}\mbox{ or }\gamma_{y}(X,R)=\frac{\Pr(Y=y,S=o\mid X,R)}{\Pr(Y=y,S=o\mid X)}.$
@@
Suppose Assumptions 1 and 2 hold.
  1. i.
    If Assumption 3(i) holds, $\pi_{d}(X,R)=\sum_{y\in\mathcal{Y}}\gamma_{y,d}(X,R)\Pr\{Y(d)=y\mid S=e,X\}\mbox{ almost surely}.$
  2. ii.
    If Assumption 3(ii) holds, $\pi_{d}(X,R)=\sum_{y\in\mathcal{Y}}\gamma_{y}(X,R)\Pr\{Y(d)=y\mid S=e,X\}\mbox{ almost surely}.$
Proof.
Consider the first statement. By (6),
$f_{R}(R\mid S=o,X,D=d,Y=y)$ $=\frac{\Pr(Y=y,D=d,S=o\mid R,X)f_{R}(R|X)}{\Pr(Y=y,D=d,S=o|X)}=\gamma_{y,d}(X,R)f_{R}(R|X)$
$f_{R}\left(R\mid D=d,S=e,X\right)$ $=\frac{\Pr(D=d,S=e\mid R,X)f_{R}(R|X)}{\Pr(D=d,S=e|X)}=\pi_{d}(X,R)f_{R}(R|X).$
Substituting these expressions into Lemma E.1(i) and canceling $f_{R}(R\mid X)$ yields the result.
Consider the second statement. By (2),
$f_{R}(R\mid Y=y,S=o,X)$ $=\frac{\Pr(Y=y,S=o\mid R,X)f_{R}(R|X)}{\Pr(Y=y,S=o|X)}=\gamma_{y}(X,R)f_{R}(R|X),$
$f_{R}\left(R\mid D=d,S=e,X\right)$ $=\frac{\Pr(D=d,S=e\mid R,X)f_{R}(R|X)}{\Pr(D=d,S=e|X)}=\pi_{d}(X,R)f_{R}(R|X).$
Substituting these expressions into Lemma E.1(ii) and canceling $f_{R}(R\mid X)$ yields the result. ∎
The previous two lemmas give the generalization of our main result, expressing the causal parameter as a conditional moment and unconditional moment. Generalizing the notation of the main text, let $\mu(d,X):=\left[\Pr\{Y(d)=y_{1}\mid S=e,X\},\ldots,\Pr\{Y(d)=y_{K-1}\mid S=e,X\}\right]^{\top}\in\mathbb{R}^{K-1}$. [^18] In particular, the main text takes $K=2$, $y_{1}=1$, and $y_{2}=0$.
Theorem E.1 (Identification as conditional moment).
Suppose Assumptions 1 and 2 hold.
  1. i.
    If Assumption 3(i) holds, then, for $d\in\{0,1\}$,
    $\mathbb{E}\left\{\widetilde{\Delta}^{e}(d,X)-\widetilde{\Delta}^{o}(d,X)^{\top}\mu(d,X)\mid X,R\right\}=0\text{ almost surely},$
    where $\widetilde{\Delta}^{e}(d,x):=\frac{1\{D=d,S=e\}}{\Pr(D=d,S=e\mid X=x)}-\frac{1\{Y=y_{K},D=d,S=o\}}{\Pr(Y=y_{K},D=d,S=o\mid X=x)}$ and
    $\widetilde{\Delta}^{o}(d,x):=\begin{bmatrix}\frac{1\{Y=y_{1},D=d,S=o\}}{\Pr(Y=y_{1},D=d,S=o\mid X=x)}-\frac{1\{Y=y_{K},D=d,S=o\}}{\Pr(Y=y_{K},D=d,S=o\mid X=x)}\\ \vdots\\ \frac{1\{Y=y_{K-1},D=d,S=o\}}{\Pr(Y=y_{K-1},D=d,S=o\mid X=x)}-\frac{1\{Y=y_{K},D=d,S=o\}}{\Pr(Y=y_{K},D=d,S=o\mid X=x)}\end{bmatrix}\in\mathbb{R}^{K-1}$
  2. ii.
    If Assumption 3(ii) holds, then, for $d\in\{0,1\}$,
    $\mathbb{E}\left\{\Delta^{e}(d,X)-\Delta^{o}(X)^{\top}\mu(d,X)\mid X,R\right\}=0\mbox{ almost surely},$
    where $\Delta^{e}(d,x):=\frac{1\{D=d,S=e\}}{\Pr(D=d,S=e\mid X=x)}-\frac{1\{Y=y_{K},S=o\}}{\Pr(Y=y_{K},S=o\mid X=x)}$ and
    $\Delta^{o}(x):=\begin{bmatrix}\frac{1\{Y=y_{1},S=o\}}{\Pr(Y=y_{1},S=o\mid X=x)}-\frac{1\{Y=y_{K},S=o\}}{\Pr(Y=y_{K},S=o\mid X=x)}\\ \vdots\\ \frac{1\{Y=y_{K-1},S=o\}}{\Pr(Y=y_{K-1},S=o\mid X=x)}-\frac{1\{Y=y_{K},S=o\}}{\Pr(Y=y_{K},S=o\mid X=x)}\end{bmatrix}\in\mathbb{R}^{K-1}.$
Proof.
We prove the second statement; the proof of the first statement is analogous. By Lemma E.2(ii), for any $r\in\mathcal{R}$ and $x\in\mathcal{X}$:
$\pi_{d}(x,r)=\sum_{j=1}^{K}\gamma_{y_{j}}(x,r)\Pr\{Y(d)=y_{j}\mid S=e,X=x\}$
where
$\pi_{d}(x,r)=\frac{\Pr(D=d,S=e\mid X=x,R=r)}{\Pr(D=d,S=e\mid X=x)}\mbox{ and }\gamma_{y_{j}}(x,r)=\frac{\Pr(Y=y_{j},S=o\mid X=x,R=r)}{\Pr(Y=y_{j},S=o\mid X=x)}.$
Since $\sum_{j=1}^{K}\Pr\{Y(d)=y_{j}\mid S=e,X=x\}=1$, we can rewrite the sum as
$\pi_{d}(x,r)$ $=\sum_{j=1}^{K-1}\gamma_{y_{j}}(x,r)\Pr\{Y(d)=y_{j}\mid S=e,X=x\}+\gamma_{y_{K}}(x,r)\left[1-\sum_{j=1}^{K-1}\Pr\{Y(d)=y_{j}\mid S=e,X=x\}\right].$
Further re-arranging then yields:
$\pi_{d}(x,r)-\gamma_{y_{K}}(x,r)=\sum_{j=1}^{K-1}\left\{\gamma_{y_{j}}(x,r)-\gamma_{y_{K}}(x,r)\right\}\Pr\{Y(d)=y_{j}\mid S=e,X=x\}.$
We then define
$\tilde{\gamma}(x,r):=\begin{bmatrix}\gamma_{y_{1}}(x,r)-\gamma_{y_{K}}(x,r)\\ \vdots\\ \gamma_{y_{K-1}}(x,r)-\gamma_{y_{K}}(x,r)\end{bmatrix},\quad\mu(d,x):=\begin{bmatrix}\Pr\{Y(d)=y_{1}\mid S=e,X=x\}\\ \vdots\\ \Pr\{Y(d)=y_{K-1}\mid S=e,X=x\}\end{bmatrix}$
and rewrite the previous display as
$\pi_{d}(x,r)-\gamma_{y_{K}}(x,r)=\tilde{\gamma}(x,r)^{\top}\mu(d,x).$
The result then follows immediately by iterated expectations. ∎
Corollary E.1 (Identification as representation).
Suppose Assumptions 1 and 2 hold. Consider any measurable function $H_{d}:\mathcal{X}\times\mathcal{R}\to\mathbb{R}^{J}$ with $J\geq K-1$.
  1. i.
    If Assumption 3(i) holds, then, for $d\in\{0,1\}$,
    $\mathbb{E}\{H_{d}(X,R)\widetilde{\Delta}^{e}(d,X)\mid X\}=\mathbb{E}\{H_{d}(X,R)\widetilde{\Delta}^{o}(d,X)^{\top}\mid X\}\mu(d,X)\mbox{ almost surely}.$
    If $\mathbb{E}\{H_{d}(X,R)\widetilde{\Delta}^{o}(d,X)^{\top}\mid X\}$ has full column rank almost surely, then $\mu(d,X)$ is identified.
  2. ii.
    If Assumption 3(ii) holds, then, for $d\in\{0,1\}$,
    $\mathbb{E}\left\{H_{d}(X,R)\Delta^{e}(d,X)\mid X\right\}=\mathbb{E}\left\{H_{d}(X,R)\Delta^{o}(X)^{\top}\mid X\right\}\mu(d,X)\mbox{ almost surely}.$
    If $\mathbb{E}\{H_{d}(X,R)\Delta^{o}(X)^{\top}\mid X\}$ has full column rank almost surely, then $\mu(d,X)$ is identified.
E.2 Choice of Representation
As in the main text, Theorem E.1 and Corollary E.1 imply that we can use existing results on conditional moment restrictions for identification and inference on the average potential outcomes in the experimental sample. Importantly, any choice of representation satisfying the full rank condition suffices for identification as well as inference on the average potential outcomes in the experimental sample, as we will discuss next.
Furthermore, we can again apply results in Chamberlain (1987) and Newey (1993) to derive the efficient representation for downstream causal inference with a discrete outcome. For exposition, consider the case in which Assumption 3(ii) holds. Theorem E.1(ii) yields the stacked conditional moment restriction:
$\mathbb{E}\left\{\begin{matrix}\Delta^{e}(0,X)-\Delta^{o}(X)^{\top}\mu(0,X)\\ \Delta^{e}(1,X)-\Delta^{o}(X)^{\top}\mu(1,X)\end{matrix}\;\Bigg|\;X,R\right\}=0.$
We take the difference of the treated and untreated moment restrictions. Generalizing the notation of the main text, let $\Delta^{e}(X)=\Delta^{e}(1,X)-\Delta^{e}(0,X)$ and $\theta(X)=\mu(1,X)-\mu(0,X)$. Then the differenced moment restriction is simply
$\mathbb{E}\left\{\Delta^{e}(X)-\Delta^{o}(X)^{\top}\theta(X)|X,R\right\}=0,$ (7)
generalizing (2) from the main text. We conclude that the efficient representation in the class of models satisfying (7) is
$H^{*}(X,R)=\frac{\mathbb{E}\{\Delta^{o}(X)\mid X,R\}}{\sigma^{2}(\theta,X,R)},\quad\sigma^{2}(\theta,X,R):=\mathbb{E}[\{\Delta^{e}(X)-\Delta^{o}(X)^{\top}\theta(X)\}^{2}\mid X,R]$
where $H^{*}(X,R)\in\mathbb{R}^{K-1}$, $\mathbb{E}\{\Delta^{o}(X)\mid X,R\}\in\mathbb{R}^{K-1}$, and $\sigma^{2}(\theta,X,R)\in\mathbb{R}$. Its $j$th component is the scalar $H_{j}^{*}(X,R)=\frac{\mathbb{E}\{\Delta_{j}^{o}(X)\mid X,R\}}{\sigma^{2}(\theta,X,R)},$ where $\Delta_{j}^{o}(X)=\frac{1\{Y=y_{j},S=o\}}{\Pr(Y=y_{j},S=o\mid X)}$.
As before, this formula reveals that the efficient representation of the RSV requires three types of predictions: (i) prediction of each outcome category given the RSV, which appears in the numerator $\mathbb{E}\{\Delta^{o}(X)\mid X,R\}$; (ii) prediction of the treatment given the RSV, which appears in the denominator $\sigma^{2}(\theta,X,R)$; and (iii) prediction of the sample indicator given the RSV, which appears in both the numerator and the denominator.
Finally, Theorem E.1 and Corollary E.1 provide testable implications of the identifying assumptions through over-identifying restrictions: we can test whether alternative choices of the representation $H(X,R)$ and $H^{\prime}(X,R)$ yield significantly different estimates of the average potential outcomes in the experimental sample.
E.3 Estimation and Inference with Discrete Covariates
In light of these identification results, we extend our inferential procedure to allow for discrete outcomes and discrete covariates. We continue to focus on the case of Assumption 3(ii) for exposition.
As in the main text, for clarity, we describe our inferential procedure using sample splitting with train and test folds; the generalization to cross-fitting is straightforward. The extension for discrete covariates is conceptually simple: for each $x\in\mathcal{X}$, we estimate the generalized conditional average treatment effect in the experimental sample $\theta(x)\in\mathbb{R}^{K-1}$. Then, we average over $X$ to estimate the generalized average treatment effect in the experimental sample $\theta\in\mathbb{R}^{K-1}$. This may be further reduced to the scalar $\theta_{0}\in\mathbb{R}$, which is the standard average treatment effect. Importantly, the inference results from Section 4 apply directly for each covariate value $x\in\mathcal{X}$.
Algorithm E.1 (Inference with discrete outcome and discrete covariates).
Given $\{S_{i},1\{S_{i}=e\}D_{i},1\{S_{i}=o\}Y_{i},X_{i},R_{i}\}$, where $Y\in\mathcal{Y}=\{y_{1},\ldots,y_{K}\}$ and $X\in\mathcal{X}=\{x_{1},\ldots,x_{|\mathcal{X}|}\}$:
@@ E.3 Estimation and Inference with Discrete Covariates
    Divide the sample into train and test folds.